# Visual Scratchpads: Enabling Global Reasoning in Vision

### v1

@@ B.1 Training procedure for the inductive scratchpad model
Consider an input image $x=f_{0}$ with scratchpad frames $f_{1},\ldots,f_{T}$ and label $y$ from the training set. The recurrent module $\mathcal{M}$ can be trained by teacher forcing, i.e., the model can be trained on samples of the type $f_{i}\to(\hat{f},\hat{y},\hat{h})=(f_{i+1},y,\mathbbm{1}(i+1=T))$. The issue with this training method is that the recurrent module $\mathcal{M}$ is solely trained on samples from the training distribution. However, during inference where scratchpad frames are not available, the recurrent module $\mathcal{M}$ will use its own generated frames as the input to itself. This discrepancy between the input distribution of the module at training and at test time could deteriorate the model’s performance. We initially implemented our model with teacher forcing training described above and observed that the model can learn all the tasks rather well. The issue, however, is that, especially at the beginning of training, the predicted frames are not guaranteed to be close enough to the train distribution to perform well during inference. Hence, we decided to use the following alternative approach. We provide a frame $f_{i}$ from the training set to the model to get the predicted next frame $\hat{f}_{i+1}$ along with the predicted label and halt variables $\hat{y}_{i+1},\hat{h}_{i+1}$. We then provide the predicted scratchpad frame to get the next frame $\hat{f}_{i+2}$ along with the next prediction for the label and halt variable $\hat{y}_{i+2},\hat{h}_{i+2}$. Finally, we compute the loss for all $\hat{f}_{i+1},\hat{y}_{i+1},\hat{h}_{i+1},\hat{f}_{i+2},\hat{y}_{i+2},\hat{h}_{i+2}$ and their corresponding values in the training set. Note that we consider $\hat{f}_{i+1}$ an independent input for the model and no gradient is backpropagated through it. As a result, during training the model’s input comes from both the training distribution and the distribution of the generated frames of the model itself. We found that this method gives a considerable increase in the training speed of our models and decided to use this method for our experiments.
@@
We note that this problem of discrepancy between training distribution and generation distribution during inference has been previously observed in settings such as text generation in recurrent neural networks (RNNs) and reinforcement learning, for instance in (Bengio et al., 2015) which proposed a scheduled sampling approach as follows: for each token, they sample it either from the train distribution with probability $\epsilon$ or from the model itself with probability $1-\epsilon$ and use a schedule (e.g., linear or exponential) to reduce $\epsilon$ during training. We note that our setting is simpler as the modeling in our setting is Markovian and each scratchpad frame is only generated based on the previous one and not the whole history in contrast to RNNs. Hence, our simplified approach of having a fixed rate of samples from the training and generation distributions worked well.
@@ Appendix C Dataset generation
For each task, we generate a dataset with 1M training samples and 10k validation samples. For both validation and training sets half of the samples have 0 and half have 1 as the label meaning that the baseline accuracy for this dataset would be $50\%$. It is important to note that these datasets are generated in a way that minimal spurious correlations are introduced, otherwise, the model might have used those correlations for weak learning and achieving better-than-random accuracies. We explain the generation algorithm for each of the datasets below.
@@ C.1 Cycles task
For the cycles task, we generate images of size $448\times 448$. We further choose the nodes randomly on an invisible circle with a radius of 220. Constraining the nodes to be on an invisible circle ensures that no three points are (almost) collinear. In this case, each node on the circle can be specified by its angle $\theta$. We also ensure that every two nodes are at least $\epsilon=0.2$ radians apart on the circle. To generate the points, we select $n-1$ random numbers between 00 and $2\pi-n\epsilon$ and then sort them: $x_{1}\leq x_{2}\leq\ldots\leq x_{n-1}$. We also select a parameter $\beta$ randomly in $[0,2\pi]$. Finally, we define the points to be
@@
Scratchpads.
For the multi-frame scratchpad of the cycles task, we first color the rightmost node in blue for the first frame. At each later frame, we color (at most) two more nodes/edges from both sides. In other words, the $k+1$th frame includes all the nodes/edges with a distance less than or equal to $k$ from the rightmost node colored in blue. Consequently, the last scratchpad frame which is the same as the single-frame scratchpad for this task colors the cycle that passes through the rightmost node in blue (whether the label is 0 or 1). We note that this resembles to what humans would naturally do by following one of the cycles (with a pen for instance).
@@ C.2 Strings task
The generation process of the strings task is similar to the cycles task. We have $2n$ invisible nodes (called anchor nodes) and these $2n$ nodes are connected with $2n$ 3rd-degree Bézier curves such that we have either two strings (label 0) or a single string (label 1) equiprobably. For this task, we also generate images of size $448\times 448$ and choose the anchor points on an invisible circle of radius $200$ with the same process described for the cycles task.
Next, we explain how Bézier curves are drawn. To specify a $k$th degree Bézier curves between points $A$ and $B$ one needs to first define $k-1$ control points $C_{1},\ldots,C_{k-1}$. To simplify the notation, we define $P_{0}=A,P_{1}=C_{1},\ldots,P_{k-1}=C_{k-1},P_{k}=B$. In this case, the Bézier curve is given by
@@
Scratchpads.
In order to generate the multi-frame scratchpad for the strings task, we use a similar procedure to what we do for the cycles task. We first color the rightmost anchor node. At each of the later frames, we extend the colored string from both sides by going to the next anchor nodes. Thus, the $k+1$th frame colors the string that passes through the rightmost anchor node up to the anchor nodes that have distance $k$ from the rightmost starting anchor node. Analogous to the cycles task, the last scratchpad frame (equivalently the single-frame scratchpad) for this task colors the string that passes through the rightmost node in blue. This is also similar to what humans would do by following one of the strings.
@@ C.3 Maze tasks
First, we explain the logic shared by both the rectangular and circular mazes. Afterward, we discuss the specifics of these two versions. Our mazes always have two parts a source/start cell colored in blue and a sink/end cell colored in red. The source and sink cell are either in one component (label 1) or not (label 0). Both rectangular and circular mazes can be viewed as graphs where each cell is a graph node and two nodes are connected if they are adjacent and there is no wall between them. We first note that each of our maze components is a tree, which ensures that all cells in one component are connected by a unique path. To generate our maze samples, we first generate a maze that is that has a single fully connected component where any two cells are connected by a unique path (the corresponding graph is a tree). Then we select the start and the end cells, and finally, we add a wall to the maze to break the maze into two components. We will next explain each of these parts in more detail.
@@
There are several algorithms for generating a maze with one component. These algorithms differ in their generation speed, the average length of the paths in the maze, and the branching factor of the maze which specifies the average number of branches in the paths of the maze. Considering these factors, we have decided to use Kruskal’s algorithm (Kruskal, 1956) for generating the mazes. Kruskal’s algorithm starts with a maze where all possible walls are drawn. Then, at each step, the algorithm selects a wall randomly and removes it if the two neighboring cells of this wall were not previously connected. This algorithm is continued until the maze is fully connected. For the start point of the maze, we select one of the cells adjacent to the first wall selected by the algorithm. We then compute the distances of all the cells to the start cell and in particular the maximum distance $d_{\max}$. Then we uniformly choose the target distance in $[d_{\max}-20,d_{\max}]$, and select the end cell such that its distance from the start cell is equal to the target distance. This approach ensures that the distance between the start and end cells is random and also large enough to make the maze challenging. Finally, we insert a wall in the maze to make two components. If the label is 0 we put this wall in the unique path that connects the start and end cells, otherwise if the label is 1 we put the wall such that the path between the start and end cells remains intact. In addition to that condition, we select the wall that minimizes the difference in the size of the two resulting components (i.e., our goal is to have components of the same size ideally).
Scratchpads.
@@
To generate the multi-frame scratchpads of the maze datasets we basically simulate a breadth-first search (BFS) from the start cell. We start from the start/source cell and for each scratchpad frame, we color any cell that is at a maximum distance of $10$ from the previously colored cells until we reach the end of the maze component or the solution is found. Note that adding cells of distance $1$ at each would have resulted in too many frames. What we do for generating the scratchpads is similar to BFS. In particular, if we define $d_{\mathrm{target}}$ equal to the distance to the end cell if they are in the same component and the maximum distance from the start cell otherwise, then the $k$th scratchpad frame colors all cells within distance $\min\{10k,d_{\mathrm{target}}\}$ from the start cell (note that we end the search once the target is reached or the whole component is explored). In this case, also, the single-frame scratchpad is the same as the final scratchpad frame in the multi-frame scratchpad.
@@
Generating OOD examples for the maze datasets is more challenging than the cycles and strings datasets since one cannot simply change the maze size as it will cause resolution inconsistencies. Thus, for the maze dataset, we use the same maze size for the training set and OOD samples. Instead, we use easier samples for training and use the normal maze task dataset described above for OOD evaluation. To generate easy samples, we choose our target distance between the start and the end cell uniformly from $[10,30]$ which is significantly smaller than $[d_{\max}-20,d_{\max}]$ used for the main dataset where $d_{\max}$ was the maximum distance from the start cell (see above). The latter ensures that the number of scratchpad frames required to solve the task when the nodes are connected is less than or equal to $3$ during training. Further, instead of trying to split the maze into two components of the same size, we try to add the wall such that the size of the component that includes the start cell is closest to $\frac{30}{d_{\max}}(\frac{\mathrm{number\;of\;cells}}{2})$. By doing the latter, we make sure that the search space seen during training (size of the component including the start cell) is smaller than the main dataset, and hence samples are easier.
@@ C.3.1 Rectangular maze specifics
Rectangular mazes are primarily specific by a number $n$ which indicates the number of rows and columns of the maze resulting in $n^{2}$ cells. E.g., maze (rect.) 32 has 1024 cells. Also, note that each cell in the rectangular maze has at most 4 neighbors.
@@ C.3.2 Circular maze specifics
Circular mazes are organized into a number of concentric rings and are primarily specified by the number of rings. The zeroth circle only includes the center of the maze and is not counted into the number of rings. The first ring contains 6 cells. For each of the next rings, the number of cells is kept fixed or is doubled. More precisely, the number of cells in the rings follows: $1,6,12,12,12,24,24,24,24,24,24,48,\ldots,48,\ldots$. Also note that the center cell in the circular maze has 6 neighbors and other cells can also have up to 5 neighbors.
@@ Appendix D Additional experiments
D.1 Additional model scaling experiments on maze (circular)

### v2

@@ B.1 Training procedure for the inductive CoS model
Consider an input image $x=f_{0}$ with CoS frames $f_{1},\ldots,f_{T}$ and label $y$ from the training set. The recurrent module $\mathcal{M}$ can be trained by teacher forcing, i.e., the model can be trained on samples of the type $f_{i}\to(\hat{f},\hat{y},\hat{h})=(f_{i+1},y,\mathbbm{1}(i+1=T))$. The issue with this training method is that the recurrent module $\mathcal{M}$ is solely trained on samples from the training distribution. However, during inference where CoS frames are not available, the recurrent module $\mathcal{M}$ will use its own generated frames as the input to itself. This discrepancy between the input distribution of the module at training and at test time could deteriorate the model’s performance. We initially implemented our model with teacher forcing training described above and observed that the model can learn all the tasks rather well. The issue, however, is that, especially at the beginning of training, the predicted frames are not guaranteed to be close enough to the training distribution to perform well during inference. Hence, we decided to use the following alternative approach. We provide a frame $f_{i}$ from the training set to the model to get the predicted next frame $\hat{f}_{i+1}$ along with the predicted label and halt variables $\hat{y}_{i+1},\hat{h}_{i+1}$. We then provide the predicted CoS frame to get the next frame $\hat{f}_{i+2}$ along with the next prediction for the label and halt variable $\hat{y}_{i+2},\hat{h}_{i+2}$. Finally, we compute the loss for all $\hat{f}_{i+1},\hat{y}_{i+1},\hat{h}_{i+1},\hat{f}_{i+2},\hat{y}_{i+2},\hat{h}_{i+2}$ and their corresponding ground truth values in the training set. Note that we consider $\hat{f}_{i+1}$ an independent input for the model and no gradient is backpropagated through it. As a result, during training the model’s input comes from both the training distribution and the distribution of the generated frames of the model itself. We found that this method gives a considerable increase in the training speed of our models and decided to use this method for our experiments.
@@
We note that this problem of discrepancy between training distribution and generation distribution during inference has been previously observed in settings such as text generation in recurrent neural networks (RNNs) and reinforcement learning, for instance in [Bengio et al., 2015] which proposed a scheduled sampling approach as follows: for each token, they sample it either from the train distribution with probability $\epsilon$ or from the model itself with probability $1-\epsilon$ and use a schedule (e.g., linear or exponential) to reduce $\epsilon$ during training. We note that our setting is simpler as the modeling in our setting is Markovian and each CoS frame is only generated based on the previous one and not the whole history in contrast to RNNs. Hence, our simplified approach of having a fixed rate of samples from the training and generation distributions worked well.
@@
B.2 Compute overhead for different methods
Taking as a reference the No-CoS baseline, we can look at the compute overhead of each proposed model. The single-frame model only adds a linear layer that is used to map the latents to the reconstructed pixels. This has a negligible impact on compute, as the backbone is a lot bigger. In terms of training time, this does not produce noticeable overheads. Similarly, the multi-frame CoS adds a linear layer for each frame of the CoS, making the compute overhead dependent on the length of the trajectory. This, together with the fact that the data loader needs to load all the target sketches for each sample, ends up causing a large increase in training time (e.g., 20x longer for maze rect. 32 dataset). On the other hand, the inductive CoS model does not need additional projection layers, therefore it does not increase the number of parameters. In the case of “CoS (only TF)” the training time is roughly the same as the single-frame model as it is trained with pairs of sketches. However, in the full inductive CoS model that also forwards generated frames, the training time doubles w.r.t. the single-frame model. At inference time, no-CoS, single-frame and multi-frame CoS have the same complexity, while the inductive CoS model scales according to the size of the problem at hand, which is a desirable property. Moreover, this overhead is completely offset by the ability to use smaller models.
@@ Appendix C Dataset generation
For each task, we generate a dataset with 1M training samples and 10k validation samples. For both validation and training sets half of the samples have 0 and half have 1 as the label meaning that the baseline accuracy for this dataset is $50\%$. It is important to note that these datasets are generated in a way that minimal spurious correlations are introduced, otherwise, the model might have used those correlations for weak learning and achieving better-than-random accuracies. We explain the generation algorithm for each of the datasets below.
@@ C.1 Cycles task
For the cycles task, we generate images of size $448\times 448$. We further choose the nodes randomly on an invisible circle with a radius of 220. Constraining the nodes to be on an invisible circle ensures that no three points are (almost) collinear. In this case, each node on the circle can be specified by its angle $\theta$. We also ensure that every two nodes are at least $\epsilon$ radians apart on the circle. To generate the points, we select $n-1$ random numbers between 00 and $2\pi-n\epsilon$ and then sort them: $x_{1}\leq x_{2}\leq\ldots\leq x_{n-1}$. We also select a parameter $\beta$ randomly in $[0,2\pi]$. Finally, we define the points to be
@@
Chain-of-sketch.
For the multi-frame CoS of the cycles task, we first color the rightmost node in blue for the first frame. At each later frame, we color (at most) two more nodes/edges from both sides. In other words, the $k+1$th frame includes all the nodes/edges with a distance less than or equal to $k$ from the rightmost node colored in blue. Consequently, the last CoS frame which is the same as the single-frame CoS for this task colors the cycle that passes through the rightmost node in blue (whether the label is 0 or 1). We note that this resembles what humans would naturally do by following one of the cycles (with a pen for instance).
@@ C.2 Strings task
The generation process of the strings task is similar to the cycles task. We have $2n$ invisible nodes (called anchor nodes) and these $2n$ nodes are connected with $2n$ 3rd-degree Bézier curves such that we have either two strings (label 0) or a single string (label 1), equiprobably. For this task, we also generate images of size $448\times 448$ and choose the anchor points on an invisible circle of radius $200$ with the same process described for the cycles task.
Next, we explain how Bézier curves are drawn. To specify a $k$th degree Bézier curve between points $A$ and $B$ one needs to first define $k-1$ control points $C_{1},\ldots,C_{k-1}$. To simplify the notation, we define $P_{0}=A,P_{1}=C_{1},\ldots,P_{k-1}=C_{k-1},P_{k}=B$. In this case, the Bézier curve is given by
@@
Chain-of-sketch.
In order to generate the multi-frame CoS for the strings task, we use a similar procedure to what we do for the cycles task. We first color the rightmost anchor node. At each of the later frames, we extend the colored string from both sides by going to the next anchor nodes. Thus, the $k+1$th frame colors the string that passes through the rightmost anchor node up to the anchor nodes that have distance $k$ from the rightmost starting anchor node. Analogous to the cycles task, the last CoS frame (equivalently the single-frame CoS) for this task colors the string that passes through the rightmost node in blue. This is also similar to what humans would do by following one of the strings.
@@ C.3 Maze tasks
First, we explain the logic shared by both the rectangular and circular mazes. Afterward, we discuss the specifics of these two versions. Our mazes always have two parts a source/start cell colored in blue and a sink/end cell colored in red. The source and sink cell are either in one component (label 1) or not (label 0). Both rectangular and circular mazes can be viewed as graphs where each cell is a graph node and two nodes are connected if they are adjacent and there is no wall between them. We first note that each of our maze components is a tree, which ensures that all cells in one component are connected by a unique path. To generate our maze samples, we first generate a maze that has a single fully connected component where any two cells are connected by a unique path (the corresponding graph is a tree). Then we select the start and the end cells, and finally, we add a wall to the maze to break the maze into two components. We will next explain each of these parts in more detail.
@@
There are several algorithms for generating a maze with one component. These algorithms differ in their generation speed, the average length of the paths in the maze, and the branching factor of the maze which specifies the average number of branches in the paths of the maze. Considering these factors, we have decided to use Kruskal’s algorithm [Kruskal, 1956] for generating the mazes. Kruskal’s algorithm starts with a maze where all possible walls are drawn. Then, at each step, the algorithm selects a wall randomly and removes it if the two neighboring cells of this wall were not previously connected. This algorithm is continued until the maze is fully connected. For the start point of the maze, we select one of the cells adjacent to the first wall selected by the algorithm. We then compute the distances of all the cells to the start cell and in particular the maximum distance $d_{\max}$. Then we uniformly choose the target distance in $[d_{\max}-20,d_{\max}]$, and select the end cell such that its distance from the start cell is equal to the target distance. This approach ensures that the distance between the start and end cells is random and also large enough to make the maze challenging. Finally, we insert a wall in the maze to make two components. If the label is 0 we put this wall in the unique path that connects the start and end cells, otherwise if the label is 1 we put the wall such that the path between the start and end cells remains intact. In addition to that condition, we select the wall that minimizes the difference in the size of the two resulting components (i.e., our goal is to have components of the same size ideally).
Chain-of-sketch.
@@
To generate the multi-frame CoS of the maze datasets we basically simulate a breadth-first search (BFS) from the start cell. We start from the start/source cell and for each CoS frame, we color any cell that is at a maximum distance of $10$ from the previously colored cells until we reach the end of the maze component or the solution is found. Note that adding cells of distance $1$ at each step would have resulted in too many frames. What we do for generating the CoS frames is similar to BFS. In particular, if we define $d_{\mathrm{target}}$ equal to the distance to the end cell if they are in the same component and the maximum distance from the start cell otherwise, then the $k$th CoS frame colors all cells within distance $\min\{10k,d_{\mathrm{target}}\}$ from the start cell (note that we end the search once the target is reached or the whole component is explored). In this case, also, the single-frame CoS is the same as the final CoS frame in the multi-frame CoS.
@@
Generating OOD examples for the maze datasets is more challenging than the cycles and strings datasets since one cannot simply change the maze size as it will cause resolution inconsistencies. Thus, for the maze dataset, we use the same maze size for the training set and OOD samples. Instead, we use easier samples for training and use the normal maze task dataset described above for OOD evaluation. To generate easy samples, we choose our target distance between the start and the end cell uniformly from $[10,30]$ which is significantly smaller than $[d_{\max}-20,d_{\max}]$ used for the main dataset where $d_{\max}$ was the maximum distance from the start cell (see above). The latter ensures that the number of CoS frames required to solve the task when the nodes are connected is less than or equal to $3$ during training. Further, instead of trying to split the maze into two components of the same size, we try to add the wall such that the size of the component that includes the start cell is closest to $\frac{30}{d_{\max}}(\frac{\mathrm{number\;of\;cells}}{2})$. By doing the latter, we make sure that the search space seen during training (size of the component including the start cell) is smaller than the main dataset, and hence samples are easier.
@@ C.3.1 Rectangular maze specifics
Rectangular mazes are primarily specified by a number $n$ which indicates the number of rows and columns of the maze resulting in $n^{2}$ cells. E.g., maze (rect.) 32 has 1024 cells. Also, note that each cell in the rectangular maze has at most 4 neighbors.
@@ C.3.2 Circular maze specifics
Circular mazes are organized into a number of concentric rings and are primarily specified by the number of rings. The zeroth circle only includes the center of the maze and is not counted in the number of rings. The first ring contains 6 cells. For each of the next rings the number of cells is kept fixed or is doubled. Also note that the center cell in the circular maze has 6 neighbors and other cells can also have up to 5 neighbors.
C.4 Image PVR tasks
In all of the experiments in this paper, we used $7\times 7$ grids. First, we randomly select one MNIST digit as the pointer from $\{1,2,3,4,5,6\}$. Next, for each of the subsequent rows, we put $k\leq 7$ CIFAR-10 images such that each image is equiprobably sampled from the airplane or car classes. $k$ is a fixed number of a specific task. The target is the result of an aggregation function applied to the row indicated by the pointer. For the main experiments of the paper, we use parity as the aggregation function, i.e., whether the number of occurrences of the first image (from left in the row) in the row is even (label 0) or odd (label 1). For experiments in Appendix D.1, we also use majority as the aggregation function where the label is 1 if the object appearing first (from left in the row) is in strict majority.
Chain-of-sketch. We only define CoS for the parity aggregation function. For chain-of-sketch, we first hide all parts of the image except the indicated row as the first image of the chain. Next at each step, we compute the parity cumulatively: we check whether the object we are checking is the same as the first object or not, if it is, we update the parity and write the updated number on that image, otherwise, we put an ’x’ on the image. We continue this until all images in the row are processed.
OOD samples. For the train distribution, we simply mix the PVR task for the number of images per row $k=2,3,4$. For the OOD evaluation set, we check the models on the task for $k=5$.
C.5 Prompts and details of experiments with multi-modal LLMs
In our experiments with GPT models, we tried various image resolutions: $256\times 256$, $512\times 512$, $1024\times 1024$, and $2048\times 2048$, as well as both high and low resolution settings of the OpenAI API. In our experiments, using high-resolution mode on $512\times 512$ images seemed to work best, and we report those results in this paper. We also experimented with different prompt formulations and found that detailed prompts tended to improve performance. The exact prompts are provided in Table 5. We also note that tool calling was disabled in our API calls and the reported results. Also we used the following checkpoints of the models: gpt-4o-2024-08-06, o4-mini-2025-04-16, and o3-2025-04-16.
Setting Prompt
Cycles with no in-context examples This image contains <NODES> nodes which are shown by large white filled dots. There are <NODES> edges between the nodes. Does the image contain a single closed cycle or two separate closed cycles? (A cycles is a closed path which passes through a number of nodes so when there is a single cycle all nodes are reachable from each other and when answer is 2 there are two separate connected components in the graph.) Provide your reasoning and end your answer with “output: 1” or “output: 2”. <task img>
Cycles with 4 in-context examples Each image contains <NODES> nodes which are shown by large white filled dots. There are <NODES> edges between the nodes. Does the fifth (last) image contain a single closed cycle or two separate closed cycles? (A cycles is a closed path which passes through a number of nodes so when there is a single cycle all nodes are reachable from each other and when answer is 2 there are two separate connected components in the graph.) For instance, in the first two images, there are two separate cycles and in the third and fourth images there is a single cycle. Provide your reasoning and end your answer with “output: 1” or “output: 2”. <img ex.1><img ex.2><img ex.3><img ex.4><task img>
Maze (rect.) with no in-context examples In the maze, is there a path between the blue square to the red square which doesn’t cross any walls (shown by thick black lines)? Try to reason. For example, you can start from the blue square and try to see whether you reach the red square or not. End your answer with a Yes or No. <task img>
Maze (rect.) with 4 in-context examples You are given 5 maze images where the task is to determine whether the blue and red square are connected or not, i.e., if there’s a path not blocked by walls (shown by thick black lines). For example, in the first two images there is no valid path between the blue and red cells. While in the third and fourth images there is a path. Now, for the fifth image, are the blue and red cells connected? Provide your reasoning and end your answer with a Yes or No. <img ex.1><img ex.2><img ex.3><img ex.4><task img>
Table 5: Prompts used for the results of Table 1. <NODES> is replaced by the number of nodes in the image. Also, <img ex.1>...<img ex.4> are the in-context image examples and <task img> is the task image.
@@ Appendix D Additional experiments
D.1 PVR experiments clarifying the concept of globality
To further clarify the notion of globality, we consider the following variant of the PVR task: instead of using the parity function on the indicated row, we use the majority function on the selected row, such that the label is 1 if the class that appears first (from left) is in the (strict) majority and 00 otherwise. Parity and majority both depend on all of the inputs. However, the majority function, in contrast to the parity function is not a global function, as even seeing one element of the majority provides non-trivial information on the output. This is while seeing all except one member of a parity function provides no information as the last member can equiprobably change the outcome. To see this better in action we check if the same PVR task in $7\times 7$ grid with 4 numbers per row and majority function aggregation is learnable or not and we compare it with the parity function. In Figure 9 we see that the majority variant of the task is learnable without CoS, whereas PVR with the parity aggregation is only learnable when (at least) a single-frame CoS is used. Here, we have conducted the experiments using the ViT-B model size, but we have also confirmed that the PVR task with parity aggregation remains hard to learn without CoS even if ViT-L or ViT-H are used.
Figure 9: PVR with majority function is learnable by the ViT-B model without CoS. Whereas, PVR with parity function is not learnable by ViT without CoS regardless of the model size.
D.2 Additional model scaling experiments on maze (circular)